OpenStudy (olivya999):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x.

OpenStudy (olivya999):

\[f(x)= \frac{ 8 }{ x }\] and \[g(x)= \frac{ 8 }{ x }\]

OpenStudy (anonymous):

holdup

OpenStudy (igreen):

Plug in \(\sf \dfrac{8}{x}\) for \(\sf x\) in \(\sf f(x)\).

OpenStudy (anonymous):

\[f(g(x))=\frac{ 8 }{ (\frac{ 8 }{ x }) }= 8 \times \frac{ x }{ 8 }\]

OpenStudy (igreen):

And vice versa.